# Mitigating Omitted Variable Bias in Empirical Software Engineering

### v2

@@ 4.5. Sensitivity Analysis with E-values
The sensitivity analysis based on simulation presented in Section 4.4 is informative and flexible, but it also has clear disadvantages: it can be very time consuming, and it does usually requires a good amount of empirical data to tune the simulation parameters in a way that is representative of the domain. As a less demanding alternative, this section presents a tipping-point analysis similar to the one we described in Section 3.2 for the other case study.
4.5.1. E-values
@@
4.5.2. Computing E-values
Let’s get into computing the E-value in our case study of team size and effort. Since we don’t have real-world data directly about the variables of Figure 7, we resort, once again, to simulation to get some plausible data in a convenient format. Unlike in Section 4.4, now we don’t need to perform many repetitions with small sample size, since the simulated data will not be used for a sensitivity analysis but to create a synthetic dataset based on the Table 4’s parameters. Thus, we simply draw ten thousand datapoints by sampling model (3) for each of the following parameter combinations:
@@ 4.5.2. Computing E-values
Figure 9. E-value tipping-point analysis for the unknown confounder $Z$ on the $T\to E$ relation in Figure 7.
@@
4.5.3. Confounding Sensitivity
9(a) plots the E-values for the three effect sizes $0.1$, $0.3$, and $0.5$. Overall, the E-value is proportional to $\delta$ (the parameter that captures the increase in treatment level caused by a possible confounder $Z$) and to $t_{e}$ (the actual effect $T\to E$). This simply reflects that the bigger the effect to be explained away by a confounder, the stronger the confounder has to sway treatment and outcome at the same time.
Then, 9(b) lists precise E-values for certain combinations of $\delta$ and $t_{e}$. Let’s look into a couple of interesting parameter combinations:
@@ 4.5.3. Confounding Sensitivity
5. Dealing with Omitted Variable Bias
@@
If the previous step’s sensitivity analysis is inconclusive, in that it failed to rule out the possibility of confounding, one can perform a more precise analysis of confounding based on simulation. Section 4.4 illustrated this on our second case study of teamwork productivity. A simulation analysis is flexible, because one can explore many different variants of generative and inference models. It also supports analyzing the impact of dealing with small sample sizes in a way that realistically reflects the availability of data in the study domain. These advantages come with a cost in terms of simulation time; usually, however, it still is much cheaper to perform a detailed simulation than to embark “blind” in running an empirical study without a clear understanding of the possible confounders, and of the threats to validity they may introduce.
@@
Proceed with caution::
@@ 6. Conclusions
First, if the causal structure among variables admits an adjustment set that only includes measurable variables, one can correct for an omitted variable bias by simply conditioning on all variables in the adjustment set. If this is not possible, one can perform a sensitivity analysis, whose goal is investigating the impact of unknown confounders. The paper presented different kinds of sensitivity analyses, including both so-called tipping-point analyses based on canonical distributional assumptions, and more precise, but also computationally expensive, analyses based on simulation. We demonstrated these techniques on two case studies—the relation between programming languages and code quality, and the effect on team size on software development effort—taken from recent statistical analyses of data in these two domains [38, 11].

### v3

@@ 4.5 Sensitivity Analysis with E-values
The sensitivity analysis based on simulation presented in Section 4.4 is informative and flexible, but it also has clear disadvantages: it can be very time consuming, and it does usually requires a good amount of empirical data to tune the simulation parameters in a way that is representative of the domain. As a less demanding alternative, this section presents a tipping-point analysis similar to the one we described in Section 3.2 for the other illustrative example.
4.5.1 E-values
@@
4.5.2 Computing E-values
Let’s get into computing the E-value in our illustrative example of team size and effort. Since we don’t have real-world data directly about the variables of Figure 7, we resort, once again, to simulation to get some plausible data in a convenient format. Unlike in Section 4.4, now we don’t need to perform many repetitions with small sample size, since the simulated data will not be used for a sensitivity analysis but to create a synthetic dataset based on the Table 4’s parameters. Thus, we simply draw ten thousand datapoints by sampling model (3) for each of the following parameter combinations:
@@ 4.5.2 Computing E-values
Figure 11: E-value tipping-point analysis for the unknown confounder $Z$ on the $T\to E$ relation in Figure 7.
@@
4.5.3 Confounding Sensitivity
11(a) plots the E-values for the three effect sizes $0.1$, $0.3$, and $0.5$. Overall, the E-value is proportional to $\delta$ (the parameter that captures the increase in treatment level caused by a possible confounder $Z$) and to $t_{e}$ (the actual effect $T\to E$). This simply reflects that the bigger the effect to be explained away by a confounder, the stronger the confounder has to sway treatment and outcome at the same time.
Then, 11(b) lists precise E-values for certain combinations of $\delta$ and $t_{e}$. Let’s look into a couple of interesting parameter combinations:
@@ 4.5.3 Confounding Sensitivity
5 Dealing with Omitted Variable Bias
@@
If the previous step’s sensitivity analysis is inconclusive, in that it failed to rule out the possibility of confounding, one can perform a more precise analysis of confounding based on simulation. Section 4.4 illustrated this on our second illustrative example of teamwork productivity. A simulation analysis is flexible, because one can explore many different variants of generative and inference models. It also supports analyzing the impact of dealing with small sample sizes in a way that realistically reflects the availability of data in the study domain. These advantages come with a cost in terms of simulation time; usually, however, it still is much cheaper to perform a detailed simulation than to embark “blind” in running an empirical study without a clear understanding of the possible confounders, and of the threats to validity they may introduce.
@@
Proceed with caution:
@@ 6 Conclusions
First, if the causal structure among variables admits an adjustment set that only includes measurable variables, one can correct for an omitted variable bias by simply conditioning on all variables in the adjustment set. If this is not possible, one can perform a sensitivity analysis, whose goal is investigating the impact of unknown confounders. The paper presented different kinds of sensitivity analyses, including both so-called tipping-point analyses based on canonical distributional assumptions, and more precise, but also computationally expensive, analyses based on simulation. We demonstrated these techniques on two illustrative examples—the relation between programming languages and code quality, and the effect on team size on software development effort—taken from recent statistical analyses of data in these two domains [60, 18].